5 over z+7= 4 over z-4
Found 3 solutions by Alan3354, stanbon, MathLover1:
Answer by Alan3354(69443)
Answer by stanbon(75887)

5 over z+7= 4 over z-4
--------------------
5/(z+7) = 4/(z-4)
-------------------------
Cross-multiply and solve for "z":
5(z-4) = 4(z+7)
5z-20 = 4z+28
z = 48
-------
Cheers,
Stan H.
==========

Answer by MathLover1(20850)